Please help me with the following problem:

Young’s rule for determining the amount of a medicine dosage for a child is given by C=ad/a+3, where "a" is the child’s age and "d" is the usual adult dosage, in milligrams. (Warning ! Do not apply this formula without checking with a physician!) An 8-year old child needs medication. What adult dosage can be used if a child’s dosage must stay between 100 mg and 200 mg?

This must be solved in the form of an inequality.
I appreciate and thank you for your help!

Answer by rfer(16322)

C=ad/a+3
200=8d/8+3
11*200=8d
2200=8d
275=d
--------------
100=8d/8+3
11*100=8d
1100=8d
137.5=d
---------------
137.5 < d < 275